15y^2-10y=5y( ) Complete the factoring.
Thank you!
oh gosh thank you jim! you're a genius!

Found 2 solutions by jim_thompson5910, whesleysgirl:
Answer by jim_thompson5910(35256)

15y^2-10y=5y(3y - 2)

Notice how 5y*3y = 15y^2 and 5y*(-2) = -10y. So this confirms the answer.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by whesleysgirl(1)